OpenStudy (anonymous):

If cos(2m) = .11, which of the following is (cos m)^2? A. .22 B. .11 C. 11.1 D. 1.11

OpenStudy (anonymous):

cos(m) 2? wut

OpenStudy (anonymous):

i fixed it @Untitled123

OpenStudy (zale101):

Do u mean 2*(cos m)^2 ?

OpenStudy (anonymous):

nope its written exactly as the question was given to me.

OpenStudy (anonymous):

@jim_thompson5910 do you understand this one?

jimthompson5910 (jim_thompson5910):

use (cos(m))^2 = ( 1 + cos(2m) )/2

OpenStudy (anonymous):

i plugged it in my calculator and i got .987948

OpenStudy (anonymous):

I got .9969 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

(cos(m))^2 = ( 1 + cos(2m) )/2 (cos(m))^2 = ( 1 + 0.11 )/2 (cos(m))^2 = ( 1.11 )/2 (cos(m))^2 = 0.555 So there has to be a typo somewhere

OpenStudy (anonymous):

i see what I did. @jim_thompson5910

OpenStudy (anonymous):

so would it be A or D? in the first part of the equation A was used but once you added it turned into D.. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

the answer is not in the choices, so like I said, there has to be a typo somewhere